FC VCTC AFC AVC ATC MCP Profit 190 100 280 250 25 112.5 98 130 260 1390 1750Fill in the chart.
Answer #1

a FC TC TR MR 0 vc 100 ... 100 100 100 190 190 190 0... 190 380 570 100 190 100 760 AFC AVC ATC MCP 100 .. ... ... ... 200 10

TC=TVC+TFC

ATC=TC/Q

AVC=VC/Q

AFC=FC/Q

MC=TCn-TCn-1